OpenStudy (anonymous):

The area of a circle is expanding at a rate of 5 inches^2/sec. At what rate is the radius changing when the radius is 8 inches.

OpenStudy (anonymous):

\[\frac{\mathrm{d}A}{\mathrm{d}t}=5 \ inch^2/s \] \[A=\pi r^2\] \[\frac{\mathrm{d}A}{\mathrm{d}t}=2\pi r \frac{\mathrm{d}r}{\mathrm{d}t} \] \[\frac{\mathrm{d}r}{\mathrm{d}t}=\frac{\mathrm{d}A}{\mathrm{d}t}\frac{1}{2 \pi r}\] then you just plug in your values for dA/dt and r ^_^ does that make sense?

OpenStudy (anonymous):

\[ \frac{dr }{ dt}= 5(\frac{ 1 }{ 2 \pi (8) })\] is this the answer

OpenStudy (anonymous):

yup yup yup ^_^